2.) Given the following information: utility function is U(x,y)=〖4x〗^0.5 y^0.5 , price of good X is N$5, the price of good Y is N$10 and the consumer income N$400.
What is the level of quantity demanded of good Y when the price of good X and Y is N$4 and N$10 respectively? Let us assume good X is on the x-axis.
3.) Given the following information: utility function is U(x,y)=〖4x〗^0.5 y^0.5 , price of good X is N$5, the price of good Y is N$10 and the consumer income N$400.
1
Expert's answer
2020-06-24T15:31:30-0400

The utility function is:


"U(x,y)=4x^{0.5} y^{0.5}"


The prices of goods x and y are "P_x = \$5\text{ and } P_y = \$10" respectively. The consumer will maximize her utility at the point where:


"\dfrac{MU_x}{MU_y} = \dfrac{P_x}{P_y}"


Therefore:


"MU_x = \dfrac{\delta U(x,y)}{\delta x} = 2x^{-0.5}y^{0.5}\\[0.3cm]
MU_y = \dfrac{\delta U(x,y)}{\delta y} = 2x^{0.5}y^{-0.5}"


Thus:


"\dfrac{2x^{-0.5}y^{0.5}}{2x^{0.5}y^{-0.5}} = \dfrac{5}{10}\\[0.3cm]
\dfrac{y}{x} = \dfrac{1}{2}\\[0.3cm]
x = 2y............................(i)"


The consumer's income is $400. Thus, the budget constraint is:


"400 = 5x + 10y"


Substituting equation (i) into the budget constraint:


"400 = 2y + 10y\\[0.3cm]
400 = 12y\\[0.3cm]
\color{red}{y^* = \dfrac{100}{3}}"


But "x = 2y" . Therefore:


"x = 2\left(\dfrac{100}{3}\right)\\[0.3cm]
\color{red}{x^* = \dfrac{200}{3}}"